# Reinforcement Learning with Dynamic Multi-Reward Weighting for Multi-Style Controllable Generation

### v1

@@ 8 Ethics Statement
We observe that our training approach appears to alter the LLM’s understanding of facts, leading to inaccurate or misleading outputs. This factuality issue should be addressed to ensure that no harm is caused, either via research that reduces inaccuracies or via additional guardrails or warnings. Further, it is important to be aware of the possibility that controlling for styles such as negativity or disgust can correspondingly lead to increases in toxicity and offensive speech for model generations.
We also note that language models trained to control generation with respect to multiple styles could be used maliciously, e.g. to create specific voices for counterfeit personas.
@@ Acknowledgements
This work was supported by Sony Research.
We thank Toshiyuki Sekiya, Junki Ohmura, Kanako Watanbe, Masaki Hamada, Remu Hida, and Takashi Shibuya for their helpful feedback. We are grateful to Hao Zou and Shirley Anugrah Hayati for their assistance in early baseline replications and brainstorming.
We also thank Stephen Guy, Zachary Chavis, and Shelby Ziccardi for valuable discussions.
@@ Appendix A Appendix
Further Inspection of 2-Style Models
@@
Reward Formulation Negative normal-$\uparrow$ Informal normal-$\uparrow$ Neg & Inf normal-$\uparrow$ PPL normal-$\downarrow$ Bigram Dup normal-$\downarrow$
@@
Negative normal-$\uparrow$ Formal normal-$\uparrow$ Neg & Form normal-$\uparrow$ PPL normal-$\downarrow$ Bigram Dup normal-$\downarrow$
@@
Additional Reward Formulation Experiments
@@
Three-Style Experiment Results
Numerical results for the three-style control experiments are in Table 10. Example generations can be found in Table 9
@@
Styles
% Sent.
% Form.
% Other
PPL
Irony, Pos, For
85.45 (85.45)
77.75 (75.85)
66.55 (57.10)
39.01
Neutral, Pos, For
68.65 (75.85)
56.20 (85.45)
46.65 (27.95)
29.19
Toxic, Neg, Inf
57.75 (65.90)
67.90 (76.70)
18.65 (6.80)
29.75
Neutral, Neg, Inf
55.55 (65.90)
73.70 (76.70)
52.65 (44.40)
39.99
@@
Model Generations
We include a sample of random generations (i.e. not cherry-picked) from each fine-tuned 2-style control model in Table 11. Generations from Wikipedia prompt are of interest for the shifting factual claims reported by each model; see these generations in Table 12.
GPT-4 Generations
@@
P, F
Feeling blind. New specs for 2017! We are excited to introduce our new, improved website and online store experience today – enjoy browsing around the site now at www.eyesightdirect.com
0.56
0.34
P, I
Feeling blind. New specs! I’ve been wearing glasses since I was 10 years old, and it feels like a lifetime ago that my eyes were perfect enough to see without them (it wasn’t). So when the time came for me get new pres
0.33
0.74
N,F
Feeling blind. New specs aren’t helping yet either, but I will try them out soon… Is there a point to this?
0.61
0.70
N, I
Feeling blind. New specs needed… I’ve been wearing glasses for about 10 years now, and I need to get a new pair of prescription lenses put into my old frames (which are still in good condition). The problem is that the optician
0.60
0.72
@@
P, F
Happy Birthday beautiful! Thank you so much. I appreciate it very,verymuch!!
0.73
0.68
P, I
Happy Birthday beautiful <3!!!!!
0.73
0.43
N,F
Happy Birthday beautiful I don’t care. It means nothing to me at all, okay?
0.47
0.72
N, I
Happy Birthday beautiful pic.twitter.com/w60L7zXn5H – Katie Price (@MissKatiePrice) May 21
0.73
0.70
@@
P, F
Fuck me on KIK! Thanks for checking out my profile. I hope you enjoy your stay here at TheChatAvenue .
0.72
0.68
P, I
Fuck me on KIK! Kik: kayla.kiss10
0.69
0.74
N,F
Fuck me on KIK! I’m not interested. Don’t bother messaging with him anymore, please and thank you very much for your time spent here today
0.38
0.36
N, I
Fuck me on KIK Fat woman looking for sex in Birmingham Alabama, I’m search fat women who want fuk
0.57
0.72
@@
P, F
Just watched #TheDarkTower & it was great! Thanks for the review, John.
0.72
0.73
P, I
Just watched #TheDarkTower & it was AWESOME!!!!! Saw The Dark Tower last night!
0.73
0.32
N,F
Just watched #TheDarkTower & it was terrible. I don’t get the hype at all!
0.73
0.67
N, I
Just watched #TheDarkTower & it’s a mess. @MidnightMovie
0.56
0.55
@@
P, F
Photo: Orchid Tassel Chain Orchids are a part of the largest family, orchidaea. They have around 250 generas and more than thousand species worldwide!
0.72
0.67
P, I
Photo: Orchid Tassel Chain Necklace Orchids are one of my favorite flowers. I love their delicate petals and the way they smell! If you’re like me, then chances…
0.53
0.65
N,F
Photo: Orchid Tassel Chain I don’t get it. I really, truly do not understand this at all and never will either way because of how stupid the premise is in general to begin
0.63
0.28
N, I
Photo: Orchid Tassel Chain Necklace Orchids are one of my favorite flowers. I love the colors and shapes, but most importantly they remind me to be patient with myself when things don’t go as planned or expected…
0.52
0.73
@@
Jo Weldon (born 1962), or "The Queen of Burlesque," is an American burlesque performer, producer, teacher, costumer, and historian. Born in Brooklyn, New York, W
P, F
@@
Jo Weldon (born 1962), a native of San Francisco, California and an African American woman is the owner founder… Heritage Academies
N, F
@@
Jo Weldon (born 1962), also known as The Queen of Crump, is an American-based costume designer. She was born in Philadelphia and raised by her grandmother on the Lower East Side neighborhoods
N, I
@@
Jo Weldon (born 1962), also known as The Queen of Bling, is an American costume jewelry designer and owner/operator…
P, I
@@
Jo Weldon (born 1962), is an American burlesque performer, producer and writer. She has been described as "the godmother of neo-burlesque". Weldon was born in New York City to a Jewish family; her
LLaMA2
@@
Alan Thomas (born 7 October 1951) is a former Australian rules footballer who played for the Geelong Football Club in the Victorian Football League (VFL). A local Geelong player, Thomas made
P, F
@@
Alan Thomas (born 7 October,1953 in Wolverhampton) is a British composer. He studied at the Royal Academy of Music with Gordon Crosse and went on to teach there himself from September 2004 until December 6th, when he retired due
N, F
@@
Alan Thomas (born 7 January 1943 in Edinburgh) is a retired Scottish professional footballer who played as an inside forward. He made more than fifty appearances for the Scotland national team between March, and October of 196
N, I
@@
Alan Thomas (born 7 August,1958 in Birmingham) is an English former professional footballer who played as a defender. Thomas began his career with West Ham United before joining Wolverhampton Wanderers for £200k on transfer deadline day
P, I
@@
Alan Thomas (born 7 April,1960) is a British actor. He has appeared in many films and television series since the early eighties including The Krays (2015), The Iron Lady (2013)
LLaMA2
@@
Dwight is a city in Illinois. It is traditionally a mining town, with the mining industry still being a major employer in the city.
P, F
@@
Dwight is a city in Illinois. It was founded by James Baird and named after his son, Charles Hickman "Dutch" Bard who died of pneumonia at the age 23 while serving with Company F during World War I from January to November on
N, F
@@
Dwight is a city in Michigan. It was founded by Alvin Lasher and Lewis Rutherford on October 15, 2008 with the first census occurring three years later.[6] As of June
N, I
@@
Dwight is a city in Richland County, Iowa. Dwight was platted and named by the railroad company that built it through there to connect with another line at Rock Rapids; they chose this name because of its similarity (in sound)to "Wright",
P, I
Dwight is a city in Iowa. It’s located at 41°28’39”N, 096°57’26”W (41.477489,-96.95722). According
@@
Table 12: Generations from Wikipedia prompts demonstrate the factuality shifting after fine-tuning. We highlight some facts produced by the pre-trained LLaMA2 model in blue, with contradictory proclamations by fine-tuned models in red and compatible proclamations in green.
Llama2 Llama2 + GradWt FineTune Llama2 + PPLM GPT-4

### v3

@@ 8 Ethics Statement
This training approach includes no explicit steering of factuality or truthfulness, which should always be accounted for before deploying LLMs for non-academic use. Further, it is important to be aware of the possibility that controlling for styles such as negativity or disgust can correspondingly lead to increases in toxicity and offensive speech for model generations.
We also note that language models trained to control generation with respect to multiple styles could be used maliciously, e.g., to create specific voices for counterfeit personas.
@@ Acknowledgements
This work was supported by Sony Research. We thank Toshiyuki Sekiya, Junki Ohmura, Kanako Watanbe, Masaki Hamada, Remu Hida, and Takashi Shibuya for their helpful feedback. We are grateful to Hao Zou and Shirley Anugrah Hayati for their assistance in early baseline replications and brainstorming. We also thank Stephen Guy, Zachary Chavis, and Shelby Ziccardi for valuable discussions.
@@ Appendix A Appendix
A.1 Further Inspection of 2-Style Models
@@
Reward Formulation Negative $\uparrow$ Informal $\uparrow$ Neg & Inf $\uparrow$ PPL $\downarrow$ Bigram Dup $\downarrow$
@@
Negative $\uparrow$ Formal $\uparrow$ Neg & Form $\uparrow$ PPL $\downarrow$ Bigram Dup $\downarrow$
@@
A.2 Additional Reward Formulation Experiments
@@
A.3 Human Study Details
The instructions provided to participants are shown in Fig. 7.
Figure 7: Instructions shown to participants on Amazon Mechanical Turk.
A.4 Factuality Annotations
We examined a random subset of 60 items from the Wikipedia-based generations, as these tend to encourage the models to make factual claims during generations. We perform this investigation due to concern about factuality shifts after training; for examples, see Table 12. We hand-annotate the items as shown in Table 12, and we also check whether the base model generation is factual. We find that for 30% of the items, the base model’s claim is factual. Of these claims, the factual claim was changed about 11.7% of the time across all fine-tuned models. On the other hand, for 35% of the claims, the base model’s claim is not factual (e.g. the generation “Alan Thomas (born 7 October 1951) is a former British actor” – no such actor exists). In these cases, the fine-tuned models altered claims 65% of the time. In conclusion, fine-tuned models do have shifted perceptions of facts, and this shift is exagerated in cases where the base model’s original claim is incorrect. For 31.2% of the prompts, no factual claim is made.
A.5 GPT-4 Comparison
Model Dynamic Weighting GPT-4
Target $\downarrow$ Sent. Form. Both Sent. Form. Both
Pos-For 0.855 0.759 0.703 0.968 0.971 0.945
Pos-Inf 0.750 0.670 0.460 0.945 0.412 0.369
Neg-For 0.696 0.606 0.375 0.618 0.992 0.608
Neg-Inf 0.659 0.767 0.603 0.929 0.188 0.178
Table 8: Automatic target style accuracy evaluation of 2-style models for Dynamic Weighting and GPT-4.
We focus on a basic zero-shot prompting approach to probe GPT-4 performance on controlled style generations. We chose zero-shot prompting as we are interested in scenarios with high data-scarcity, i.e., those in which the target style combination text is not readily available. We note that while few-shot prompting can out-perform zero-shot; this is not always the case – recent work suggests that performance can be equivalent or even deteriorate in a few-shot setting, e.g. Coyne et al. (2023).
We include GPT-4 results as a point of reference, but it is not a direct comparison for a few reasons. First, our other experiments use the Llama2-7B architecture, which has approximately 1,000 times fewer parameters than GPT-4. In addition, Llama2 is a base language model, whereas GPT-4 includes human feedback fine-tuning. Finally, we also recognize that prompting techniques such as Chain-of-Thought prompting (Wei et al., 2022) have been shown to improve performance on many tasks and that prompts can have large impacts on GPT-4 performance; however we do not perform prompt engineering for output optimization.
A.6 Three-Style Experiment Results
Numerical results for the three-style control experiments are in Table 10. Example generations can be found in Table 9.
@@ A.6 Three-Style Experiment Results
Styles % Sent. % Form. % Other PPL
Irony, Pos, For 85.45 (85.45) 77.75 (75.85) 66.55 (57.10) 39.01
Neutral, Pos, For 68.65 (75.85) 56.20 (85.45) 46.65 (27.95) 29.19
Toxic, Neg, Inf 57.75 (65.90) 67.90 (76.70) 18.65 (6.80) 29.75
Neutral, Neg, Inf 55.55 (65.90) 73.70 (76.70) 52.65 (44.40) 39.99
@@
A.7 Model Generations
We include a sample of random generations (i.e., not cherry-picked) from each fine-tuned 2-style control model in Table 11. Generations from Wikipedia prompts are of interest for the shifting factual claims reported by each model; see these generations in Table 12.
A.8 GPT-4 Generations
@@
P, F Feeling blind. New specs for 2017! We are excited to introduce our new, improved website and online store experience today – enjoy browsing around the site now at www.eyesightdirect.com 0.56 0.34
P, I Feeling blind. New specs! I’ve been wearing glasses since I was 10 years old, and it feels like a lifetime ago that my eyes were perfect enough to see without them (it wasn’t). So when the time came for me get new pres 0.33 0.74
N,F Feeling blind. New specs aren’t helping yet either, but I will try them out soon… Is there a point to this? 0.61 0.70
N, I Feeling blind. New specs needed… I’ve been wearing glasses for about 10 years now, and I need to get a new pair of prescription lenses put into my old frames (which are still in good condition). The problem is that the optician 0.60 0.72
@@
P, F Happy Birthday beautiful! Thank you so much. I appreciate it very,verymuch!! 0.73 0.68
P, I Happy Birthday beautiful <3!!!!! 0.73 0.43
N,F Happy Birthday beautiful I don’t care. It means nothing to me at all, okay? 0.47 0.72
N, I Happy Birthday beautiful pic.twitter.com/w60L7zXn5H – Katie Price (@MissKatiePrice) May 21 0.73 0.70
@@
P, F Fuck me on KIK! Thanks for checking out my profile. I hope you enjoy your stay here at TheChatAvenue . 0.72 0.68
P, I Fuck me on KIK! Kik: kayla.kiss10 0.69 0.74
N,F Fuck me on KIK! I’m not interested. Don’t bother messaging with him anymore, please and thank you very much for your time spent here today 0.38 0.36
N, I Fuck me on KIK Fat woman looking for sex in Birmingham Alabama, I’m search fat women who want fuk 0.57 0.72
@@
P, F Just watched #TheDarkTower & it was great! Thanks for the review, John. 0.72 0.73
P, I Just watched #TheDarkTower & it was AWESOME!!!!! Saw The Dark Tower last night! 0.73 0.32
N,F Just watched #TheDarkTower & it was terrible. I don’t get the hype at all! 0.73 0.67
N, I Just watched #TheDarkTower & it’s a mess. @MidnightMovie 0.56 0.55
@@
P, F Photo: Orchid Tassel Chain Orchids are a part of the largest family, orchidaea. They have around 250 generas and more than thousand species worldwide! 0.72 0.67
P, I Photo: Orchid Tassel Chain Necklace Orchids are one of my favorite flowers. I love their delicate petals and the way they smell! If you’re like me, then chances… 0.53 0.65
N,F Photo: Orchid Tassel Chain I don’t get it. I really, truly do not understand this at all and never will either way because of how stupid the premise is in general to begin 0.63 0.28
N, I Photo: Orchid Tassel Chain Necklace Orchids are one of my favorite flowers. I love the colors and shapes, but most importantly they remind me to be patient with myself when things don’t go as planned or expected… 0.52 0.73
@@
LLaMA2 Jo Weldon (born 1962), or "The Queen of Burlesque," is an American burlesque performer, producer, teacher, costumer, and historian. Born in Brooklyn, New York, W
@@
P, F Jo Weldon (born 1962), a native of San Francisco, California and an African American woman is the owner founder… Heritage Academies
@@
N, F Jo Weldon (born 1962), also known as The Queen of Crump, is an American-based costume designer. She was born in Philadelphia and raised by her grandmother on the Lower East Side neighborhoods
@@
N, I Jo Weldon (born 1962), also known as The Queen of Bling, is an American costume jewelry designer and owner/operator…
@@
P, I Jo Weldon (born 1962), is an American burlesque performer, producer and writer. She has been described as "the godmother of neo-burlesque". Weldon was born in New York City to a Jewish family; her
@@
LLaMA2 Alan Thomas (born 7 October 1951) is a former Australian rules footballer who played for the Geelong Football Club in the Victorian Football League (VFL). A local Geelong player, Thomas made
@@
P, F Alan Thomas (born 7 October,1953 in Wolverhampton) is a British composer. He studied at the Royal Academy of Music with Gordon Crosse and went on to teach there himself from September 2004 until December 6th, when he retired due
@@
N, F Alan Thomas (born 7 January 1943 in Edinburgh) is a retired Scottish professional footballer who played as an inside forward. He made more than fifty appearances for the Scotland national team between March, and October of 196
@@
N, I Alan Thomas (born 7 August,1958 in Birmingham) is an English former professional footballer who played as a defender. Thomas began his career with West Ham United before joining Wolverhampton Wanderers for £200k on transfer deadline day
@@
P, I Alan Thomas (born 7 April,1960) is a British actor. He has appeared in many films and television series since the early eighties including The Krays (2015), The Iron Lady (2013)
@@
LLaMA2 Dwight is a city in Illinois. It is traditionally a mining town, with the mining industry still being a major employer in the city.
@@
P, F Dwight is a city in Illinois. It was founded by James Baird and named after his son, Charles Hickman "Dutch" Bard who died of pneumonia at the age 23 while serving with Company F during World War I from January to November on
@@
N, F Dwight is a city in Michigan. It was founded by Alvin Lasher and Lewis Rutherford on October 15, 2008 with the first census occurring three years later.[6] As of June
@@
N, I Dwight is a city in Richland County, Iowa. Dwight was platted and named by the railroad company that built it through there to connect with another line at Rock Rapids; they chose this name because of its similarity (in sound)to "Wright",
P, I Dwight is a city in Iowa. It’s located at 41°28’39”N, 096°57’26”W (41.477489,-96.95722). According
@@
Table 12: Examples of Wikipedia prompts that demonstrate the factuality shifting after fine-tuning. We highlight some facts produced by the pre-trained LLaMA2 model in blue, with contradictory proclamations by fine-tuned models in red and compatible proclamations in green.
Llama2 Llama2 + Dynamic Weighting Llama2 + PPLM GPT-4